# The Landscape of Agentic Reinforcement Learning for LLMs: A Survey

### v1

@@ 4.8 Other Tasks
ARIA [444] compresses the sprawling action space via intention-driven reward aggregation, reducing sparsity and variance. GiGPO [110] enhances temporal credit assignment through hierarchical grouping without added computational burden. RAGEN [445] ensures stable multi-turn learning by filtering trajectories and stabilizing gradients, while advocating for reasoning-aware rewards. SPA-RL [111] decomposes delayed rewards into per-step signals, improving performance and grounding accuracy. Trinity-RFT [446] provides a unified, modular framework for reinforcement fine-tuning across tasks—including text games—enabling flexible, efficient, and scalable experimentation with diverse RL modes and data pipelines.
@@
SkyRL-SQL [447] introduces a data-efficient, multi-turn RL pipeline for Text-to-SQL, enabling LLM agents to interactively probe databases, refine, and verify SQL queries. With just 653 training examples, the SkyRL-SQL-7B model surpasses both GPT-4o and o4-mini on SQL generation benchmarks. MSRL [448] introduces multimodal structured reinforcement learning with multi-granularity rewards to overcome the SFT plateau in chart-to-code generation, achieving state-of-the-art performance on chart understanding benchmarks
@@
Time-R1 [449] enhances moderate-sized LLMs with comprehensive temporal reasoning abilities through a progressive reinforcement learning curriculum and a dynamic rule-based reward system. TimeMaster [450] trains time-series MLLMs that combine SFT with GRPO to enable structured, interpretable temporal reasoning over visualized time-series inputs.
@@
Agent models [451] internalize chain-of-action generation to enable autonomous and efficient decision-making through a combination of supervised fine-tuning and reinforcement learning. L-Zero [452] enables large language models to become general-purpose agents through a scalable, end-to-end reinforcement learning pipeline utilizing a low-cost, extensible, and sandboxed concurrent agent worker pool.
@@
Sotopia-RL [453] refines coarse episode-level rewards into utterance-level, multi-dimensional signals to enable efficient and stable RL training for socially intelligent LLMs under partial observability and multi-faceted objectives. [454] introduces an Adaptive Mode Learning (AML) framework with the Adaptive Mode Policy Optimization (AMPO) algorithm, which uses reinforcement learning to dynamically switch between multi-granular reasoning modes in social intelligence tasks, achieving higher accuracy and shorter reasoning chains than fixed-depth RL methods like GRPO.
@@ 5.1 Environment Simulator
[…]
SecRepoBench [477] ①, ③ Coding, Security Text -
[…]
@@ 5.1.1 Web Environments
In the realm of web-based environments, several benchmarks offer controlled yet realistic static environments for Agentic RL. WebShop [464] is a simulated e-commerce website featuring a large catalog of real-world products and crowd-sourced text instructions. Agents navigate various webpage types and issue diverse actions (e.g., searching, selecting items, customizing, purchasing) to find and buy products, with its deterministic search engine aiding reproducibility. Furthermore, Mind2Web [465] is a dataset designed for generalist web agents, featuring a substantial number of tasks from many real-world websites across diverse domains. It provides webpage snapshots and crowdsourced action sequences for tasks like finding flights or interacting with social profiles, emphasizing generalization across unseen websites and domains. Similarly, WebArena [466] and its multimodal extension, VisualwebArena [467], are self-hostable, reproducible web environments delivered as Docker containers. WebArena features fully functional websites across common domains like e-commerce, social forums, collaborative development, and content management systems, enriched with utility tools and knowledge bases, and supports multi-tab tasks and user role simulation. VisualwebArena extends this by introducing new tasks requiring visual comprehension and a “Set-of-Marks” (SoM) representation to annotate interactable elements on screenshots, bridging the gap for multimodal web agents. Additionally, AppWorld [469] constitutes an environment simulating a multi-application ecosystem, encompassing 9 daily-use applications (e.g., Amazon, Spotify, Gmail) with 457 invokable APIs, and constructing a digital world featuring approximately 100 virtual characters and their social relationships. Agents accomplish complex tasks (such as travel planning and social relationship management) by writing code to call APIs. In these environments, all changes to the web pages or visual elements occur exclusively in response to the agent’s actions.
@@ 5.1.2 GUI Environments
AndroidWorld [470] exemplifies such dynamism as a benchmarking environment operating on a live Android emulator, featuring 116 hand-crafted tasks across 20 real-world applications. Its dynamic nature is underscored by parameter instantiation that generates millions of unique task variations, ensuring the environment evolves into novel configurations without direct agent influence. Agents interact through a consistent interface (supporting screen interactions, app navigation, and text input) while receiving real-time state feedback, with integration to MiniWoB++ providing durable reward signals for evaluating adaptive performance. OSWorld [471] is a scalable real computer environment for multimodal agents, supporting task setup and execution-based evaluation across Ubuntu, Windows, and macOS. It includes a substantial number of real-world computer tasks involving real web and desktop applications, OS file I/O, and workflows spanning multiple applications, where all OS state changes are exclusively triggered by the agent’s actions.
@@ 5.1.3 Coding & Software Engineering Environments
Several environments instantiate agent–environment interaction under software engineering workflows. Debug-Gym [472] is a text-based interactive coding environment for LLM agents in debugging settings. It equips agents with tools like a Python debugger (pdb) to actively explore and modify buggy codebases, supporting repository-level information handling and ensuring safety via Docker containers. R2E-Gym [478] constructs a procedurally generated, executable gym-style environment of over 8K software engineering tasks, powered by the SWE-Gen pipeline and hybrid verifiers. TheAgentCompany [475] simulates a software development company, where agents act as "digital workers" performing professional tasks such as web browsing, coding, program execution, and communication with simulated colleagues. It features a diverse set of long-horizon tasks with checkpoints for partial credit, providing a comprehensive testbed for agents in a realistic workplace setting. In all these environments, the underlying problem definitions and codebases remain fixed, and changes occur solely as a result of the agent’s actions.
@@
A wide range of benchmarks and datasets focus on constructing curated task suites and evaluation pipelines. HumanEval [479] introduces a benchmark of 164 hand-crafted Python programming tasks to measure functional correctness via the pass@k metric. MBPP  [480] provides 974 entry-level Python tasks with natural language descriptions for evaluating short program synthesis. BigCodeBench [481] proposes a large-scale, contamination-free function-level benchmark of 1,140 tasks requiring composition of multiple function calls. LiveCodeBench [482] builds a continuously updated, contamination-free benchmark from real competition problems. SWE-bench [483] introduces a dynamic, execution-driven code repair benchmark derived from real GitHub issues. SWE-rebench [484] introduces a continual GitHub-mining pipeline (>>21k tasks) for both training and evaluation. DevBench [485] evaluates end-to-end development across design, setup, implementation, and testing. ProjectEval [486] constructs LLM-generated, human-reviewed project tasks with simulated user interactions. ColBench [145] instantiates multi-turn backend/frontend tasks with a privileged critic for step-wise rewards. NoCode-bench [488] evaluates LLMs on feature addition from documentation updates across real codebases. CodeBoost [279] serves as a data-centric, execution-driven training pipeline by extracting and augmenting code snippets.
@@ 5.1.4 Domain-specific Environments
ScienceWorld [458] integrates science simulations (e.g., thermodynamics, electricity, chemistry) into complex text-based tasks designed around elementary-level science education. PaperBench [490] evaluates the ability of LLM agents to replicate cutting-edge machine learning research by reproducing 20 ICML 2024 papers from scratch, scored against rubric-based subtasks. $\tau$-bench [474] simulates dynamic conversations for software engineering tasks, operating with an underlying database state and domain-specific rules that change only through the agent’s API calls.
@@
MLE-Dojo [473] is a Gym-style framework for iterative machine learning engineering workflows, built upon real-world Kaggle competitions. It provides an interactive environment for agents to iteratively experiment, debug, and refine solutions. MLE-Bench [489] establishes a benchmark for MLE by curating 75 Kaggle competitions, evaluating agents against human baselines on public leaderboards. DA-Code [487] addresses agentic data-science workflows grounded in real datasets and executable analysis, providing a focused benchmark for this domain.
@@
MedAgentGym [476] provides a domain-specific environment for biomedical code generation and testing, focusing on tasks within this specialized scientific field.
@@
SecRepoBench [477] is a domain-specific benchmark for security vulnerability repair, covering 27 repositories and 15 Common Weakness Enumeration (CWE) categories.
@@ 5.1.5 Simulated & Game Environments
Text-based environments simulate interactive settings where agent actions are expressed through natural language. LMRL-Gym [455] provides a benchmark for evaluating reinforcement learning algorithms in multi-turn language interactions, including tasks like “20 Questions” and Chess. TextWorld [457] is a sandbox environment for training agents in text-based games, offering both hand-authored and procedurally generated games. Game-based environments also emphasize visual settings that may evolve independently. Crafter [491] is a 2D open-world survival game that benchmarks deep exploration and long-horizon reasoning. Craftax [492], built upon Crafter using JAX, introduces increased complexity and GPU-acceleration for open-ended RL. The modified Crafter variant by ELLM [493] expands the action space and introduces distractor tasks. For multi-agent coordination, SMAC [494] provides StarCraft II-based benchmarks for cooperative decentralized control. Factorio [495] presents a dynamic, tick-based industrial simulation where agent inaction still alters the world state.
@@ 5.1.6 General-Purpose Environments
Some environments and benchmarks are designed for broad evaluation or to improve general agent capabilities. AgentGym [459] focuses on improving LLM agent generalization via instruction tuning and self-correction, operating on deterministic environments such as AlfWorld, BabyAI, and SciWorld. Agentbench [460] serves as a broad evaluation framework, assessing LLMs as agents across a variety of distinct interactive environments, including SQL-based, game-based, and web-based scenarios. InternBootcamp [461] is a scalable framework integrating over 1000 verifiable reasoning tasks, spanning programming, logic puzzles, and games, with a standardized interface for RL training and automated task generation.

### v4

@@ 4.8 Other Tasks
ARIA (Yang et al., 2025e) compresses the sprawling action space via intention-driven reward aggregation, reducing sparsity and variance. GiGPO (Feng et al., 2025b) enhances temporal credit assignment through hierarchical grouping without added computational burden. RAGEN (Wang et al., 2025v) ensures stable multi-turn learning by filtering trajectories and stabilizing gradients, while advocating for reasoning-aware rewards. SPA-RL (Wang et al., 2025b) decomposes delayed rewards into per-step signals, improving performance and grounding accuracy. Trinity-RFT (Pan et al., 2025) provides a unified, modular framework for reinforcement fine-tuning across tasks—including text games—enabling flexible, efficient, and scalable experimentation with diverse RL modes and data pipelines.
@@
SkyRL-SQL (Liu et al., 2025j) introduces a data-efficient, multi-turn RL pipeline for Text-to-SQL, enabling LLM agents to interactively probe databases, refine, and verify SQL queries. With just 653 training examples, the SkyRL-SQL-7B model surpasses both GPT-4o and o4-mini on SQL generation benchmarks. MSRL (Chen et al., 2025a) introduces multimodal structured reinforcement learning with multi-granularity rewards to overcome the SFT plateau in chart-to-code generation, achieving state-of-the-art performance on chart understanding benchmarks.
@@
Time-R1 (Liu et al., 2025t) enhances moderate-sized LLMs with comprehensive temporal reasoning abilities through a progressive reinforcement learning curriculum and a dynamic rule-based reward system. TimeMaster (Zhang et al., 2025m) trains time-series MLLMs that combine SFT with GRPO to enable structured, interpretable temporal reasoning over visualized time-series inputs.
@@
Agent models (Zhang et al., 2025w) internalize chain-of-action generation to enable autonomous and efficient decision-making through a combination of supervised fine-tuning and reinforcement learning. L-Zero (Zhang et al., 2025l) enables large language models to become general-purpose agents through a scalable, end-to-end reinforcement learning pipeline utilizing a low-cost, extensible, and sandboxed concurrent agent worker pool.
@@
Sotopia-RL (Yu et al., 2025c) refines coarse episode-level rewards into utterance-level, multi-dimensional signals to enable efficient and stable RL training for socially intelligent LLMs under partial observability and multi-faceted objectives. Wang et al. (2025h) introduces an Adaptive Mode Learning (AML) framework with the Adaptive Mode Policy Optimization (AMPO) algorithm, which uses reinforcement learning to dynamically switch between multi-granular reasoning modes in social intelligence tasks, achieving higher accuracy and shorter reasoning chains than fixed-depth RL methods like GRPO.
Agentic Capability
Application Planning Tool-Use Memory Self-Imp. Reasoning Percep.
Search $\bullet$ $\bullet$ $\circ$ $\circ$ $\bullet$
Code $\circ$ $\bullet$ $\circ$ $\circ$ $\bullet$
Math $\bullet$ $\circ$ $\circ$ $\bullet$
GUI $\bullet$ $\bullet$ $\circ$ $\circ$ $\bullet$
Vision $\circ$ $\circ$ $\bullet$ $\bullet$
Embodied $\bullet$ $\circ$ $\bullet$ $\circ$ $\bullet$
MAS $\bullet$ $\circ$ $\circ$ $\bullet$ $\circ$
Table 9: Application-capability dependency matrix. Dots indicate qualitative dependency levels: $\bullet$ Core, $\circ$ Supporting, – Minimal. The heatmap provides a navigation aid linking the capability taxonomy (Section 3) with the application domains (Section 6).
@@ 5.1 Environment Simulator
LMRL-Gym (Abdulhai et al., 2025) ①, ④ Interaction Text GitHub
ALFWorld (Shridhar et al., 2021) ②, ① Embodied, Text Games Text GitHub Website
TextWorld (Côté et al., 2019) ②, ① Text Games Text GitHub
ScienceWorld (Wang et al., 2022) ①, ② Embodied, Science Text GitHub Website
AgentGym (Xi et al., 2025) ①, ④ Text Games Text GitHub Website
Agentbench (Liu et al., 2024b) General Text, Visual GitHub
InternBootcamp (Li et al., 2025g) General, Coding, Logic Text GitHub
LoCoMo (Maharana et al., 2024) Interaction Text GitHub Website
MemoryAgentBench (Hu et al., 2025e) Interaction Text GitHub
WebShop (Yao et al., 2022) ②, ③ Web Text GitHub Website
Mind2Web (Gou et al., 2025) ②, ③ Web Text, Visual GitHub Website
WebArena (Zhou et al., 2024b) ②, ③ Web Text GitHub Website
VisualwebArena (Koh et al., 2024) ①, ②, ③ Web Text, Visual GitHub Website
AppBench (Wang et al., 2024b) ②, ③ App Text GitHub
AppWorld (Trivedi et al., 2024) ②, ③ App Text GitHub Website
AndroidWorld (Rawles et al., 2025) ②, ③ GUI, App Text, Visual GitHub
OSWorld (Xie et al., 2024) ②, ③ GUI, OS Text, Visual GitHub Website
WindowsAgentArena (Bonatti et al., 2024) GUI, OS Text, Visual GitHub Website
Debug-Gym (Yuan et al., 2025c) ①, ③ SWE Text GitHub Website
MLE-Dojo (Qiang et al., 2025) ②, ① MLE Text GitHub Website
$\tau$-bench (Barres et al., 2025) ①, ③ SWE Text GitHub
TheAgentCompany (Xu et al., 2024a) ②, ③, ⑤ SWE Text GitHub Website
MedAgentGym (Xu et al., 2025c) Science Text GitHub
SecRepoBench (Dilgren et al., 2025) ①, ③ Coding, Security Text -
R2E-Gym (Jain et al., 2025c) ①, ② SWE Text GitHub Website
BigCodeBench (Zhuo et al., 2025) Coding Text GitHub Website
LiveCodeBench (Jain et al., 2025b) Coding Text GitHub Website
SWE-bench (Jimenez et al., 2024) ①, ③ SWE Text GitHub Website
SWE-rebench (Badertdinov et al., 2025) ①, ③ SWE Text Website
DevBench (Li et al., 2025a) ②, ① SWE Text GitHub
ProjectEval (Liu et al., 2025g) ②, ① SWE Text GitHub Website
DA-Code (Huang et al., 2024b) ①, ③ Data Science, SWE Text GitHub Website
ColBench (Zhou et al., 2025e) ②, ① SWE, Web Dev Text GitHub Website
NoCode-bench (Deng et al., 2025a) ②, ① SWE Text GitHub Website
MLE-Bench (Chan et al., 2025) ②, ①, ③ MLE Text GitHub Website
PaperBench (Starace et al., 2025) ②, ①, ③ MLE Text GitHub Website
Crafter (Hafner, 2022) ②, ④ Game Visual GitHub Website
Craftax (Matthews et al., 2024) ②, ④ Game Visual GitHub
ELLM (Crafter variant) (Du et al., 2023) ②, ① Game Visual GitHub Website
SMAC / SMAC-Exp (Samvelyan et al., 2019) ⑤, ② Game Visual GitHub
Factorio (Hopkins et al., 2025) ②, ① Game Visual GitHub Website
SMAC-Hard (Deng et al., 2024b) ②, ④ Game Visual GitHub
TacticCraft (Ma et al., 2025a) ②, ⑤ Game Text -
@@ 5.1.1 Web Environments
In the realm of web-based environments, several benchmarks offer controlled yet realistic static environments for Agentic RL. WebShop (Yao et al., 2022) is a simulated e-commerce website featuring a large catalog of real-world products and crowdsourced text instructions. Agents navigate various webpage types and issue diverse actions (e.g., searching, selecting items, customizing, purchasing) to find and buy products, with its deterministic search engine aiding reproducibility. Furthermore, Mind2Web (Gou et al., 2025) is a dataset designed for generalist web agents, featuring a substantial number of tasks from many real-world websites across diverse domains. It provides webpage snapshots and crowdsourced action sequences for tasks like finding flights or interacting with social profiles, emphasizing generalization across unseen websites and domains. Similarly, WebArena (Zhou et al., 2024b) and its multimodal extension, VisualwebArena (Koh et al., 2024), are self-hostable, reproducible web environments delivered as Docker containers. WebArena features fully functional websites across common domains like e-commerce, social forums, collaborative development, and content management systems, enriched with utility tools and knowledge bases, and supports multi-tab tasks and user role simulation. VisualwebArena extends this by introducing new tasks requiring visual comprehension and a “Set-of-Marks” (SoM) representation to annotate interactable elements on screenshots, bridging the gap for multimodal web agents. Additionally, AppWorld (Trivedi et al., 2024) constitutes an environment simulating a multi-application ecosystem, encompassing 9 daily-use applications (e.g., Amazon, Spotify, Gmail) with 457 invokable APIs, and constructing a digital world featuring approximately 100 virtual characters and their social relationships. Agents accomplish complex tasks (such as travel planning and social relationship management) by writing code to call APIs. In these environments, all changes to the web pages or visual elements occur exclusively in response to the agent’s actions.
@@ 5.1.2 GUI Environments
AndroidWorld (Rawles et al., 2025) exemplifies such dynamism as a benchmarking environment operating on a live Android emulator, featuring 116 hand-crafted tasks across 20 real-world applications. Its dynamic nature is underscored by parameter instantiation that generates millions of unique task variations, ensuring the environment evolves into novel configurations without direct agent influence. Agents interact through a consistent interface (supporting screen interactions, app navigation, and text input) while receiving real-time state feedback, with integration to MiniWoB++ providing durable reward signals for evaluating adaptive performance. OSWorld (Xie et al., 2024) is a scalable real computer environment for multimodal agents, supporting task setup and execution-based evaluation across Ubuntu, Windows, and macOS. It includes a substantial number of real-world computer tasks involving real web and desktop applications, OS file I/O, and workflows spanning multiple applications, where all OS state changes are exclusively triggered by the agent’s actions.
@@ 5.1.3 Coding & Software Engineering Environments
Several environments instantiate agent–environment interaction under software engineering workflows. Debug-Gym (Yuan et al., 2025c) is a text-based interactive coding environment for LLM agents in debugging settings. It equips agents with tools like a Python debugger (pdb) to actively explore and modify buggy codebases, supporting repository-level information handling and ensuring safety via Docker containers. R2E-Gym (Jain et al., 2025c) constructs a procedurally generated, executable gym-style environment of over 8K software engineering tasks, powered by the SWE-Gen pipeline and hybrid verifiers. TheAgentCompany (Xu et al., 2024a) simulates a software development company, where agents act as "digital workers" performing professional tasks such as web browsing, coding, program execution, and communication with simulated colleagues. It features a diverse set of long-horizon tasks with checkpoints for partial credit, providing a comprehensive testbed for agents in a realistic workplace setting. In all these environments, the underlying problem definitions and codebases remain fixed, and changes occur solely as a result of the agent’s actions.
@@
A wide range of benchmarks and datasets focus on constructing curated task suites and evaluation pipelines. HumanEval (Chen et al., 2021) introduces a benchmark of 164 hand-crafted Python programming tasks to measure functional correctness via the pass@k metric. MBPP (Austin et al., 2021) provides 974 entry-level Python tasks with natural language descriptions for evaluating short program synthesis. BigCodeBench (Zhuo et al., 2025) proposes a large-scale, contamination-free function-level benchmark of 1,140 tasks requiring composition of multiple function calls. LiveCodeBench (Jain et al., 2025b) builds a continuously updated, contamination-free benchmark from real competition problems. SWE-bench (Jimenez et al., 2024) introduces a dynamic, execution-driven code repair benchmark derived from real GitHub issues. SWE-rebench (Badertdinov et al., 2025) introduces a continual GitHub-mining pipeline (>>21k tasks) for both training and evaluation. DevBench (Li et al., 2025a) evaluates end-to-end development across design, setup, implementation, and testing. ProjectEval (Liu et al., 2025g) constructs LLM-generated, human-reviewed project tasks with simulated user interactions. ColBench (Zhou et al., 2025e) instantiates multi-turn backend/frontend tasks with a privileged critic for step-wise rewards. NoCode-bench (Deng et al., 2025a) evaluates LLMs on feature addition from documentation updates across real codebases. CodeBoost (Wang et al., 2025n) serves as a data-centric, execution-driven training pipeline by extracting and augmenting code snippets.
Programmatic World-Model Environments.
Beyond isolated coding tasks, recent benchmarks evaluate whether agents can induce executable world models. The Code World Models Benchmark (CWMB) (Dainese et al., 2024) requires agents to synthesize Python “Environment” classes (specifically the “step” function) to replicate ground-truth dynamics, assessing both transition fidelity and downstream planning utility. Complementing this, the Code Simulation suite (Malfa et al., 2024; 2025) offers finer-grained tests on line-by-line execution prediction and algorithmic generalization. Collectively, these tasks shift the evaluation focus from functional correctness to the dynamics-induction and program-simulation capabilities essential for constructing programmatic world models.
@@ 5.1.4 Domain-specific Environments
ScienceWorld (Wang et al., 2022) integrates science simulations (e.g., thermodynamics, electricity, chemistry) into complex text-based tasks designed around elementary-level science education. PaperBench (Starace et al., 2025) evaluates the ability of LLM agents to replicate cutting-edge machine learning research by reproducing 20 ICML 2024 papers from scratch, scored against rubric-based subtasks. $\tau$-bench (Barres et al., 2025) simulates dynamic conversations for software engineering tasks, operating with an underlying database state and domain-specific rules that change only through the agent’s API calls.
@@
MLE-Dojo (Qiang et al., 2025) is a Gym-style framework for iterative machine learning engineering workflows, built upon real-world Kaggle competitions. It provides an interactive environment for agents to iteratively experiment, debug, and refine solutions. MLE-Bench (Chan et al., 2025) establishes a benchmark for MLE by curating 75 Kaggle competitions, evaluating agents against human baselines on public leaderboards. DA-Code (Huang et al., 2024b) addresses agentic data-science workflows grounded in real datasets and executable analysis, providing a focused benchmark for this domain.
@@
MedAgentGym (Xu et al., 2025c) provides a domain-specific environment for biomedical code generation and testing, focusing on tasks within this specialized scientific field.
@@
SecRepoBench (Dilgren et al., 2025) is a domain-specific benchmark for security vulnerability repair, covering 27 repositories and 15 Common Weakness Enumeration (CWE) categories.
@@ 5.1.5 Simulated & Game Environments
Text-based environments simulate interactive settings where agent actions are expressed through natural language. LMRL-Gym (Abdulhai et al., 2025) provides a benchmark for evaluating reinforcement learning algorithms in multi-turn language interactions, including tasks like “20 Questions” and Chess. TextWorld (Côté et al., 2019) is a sandbox environment for training agents in text-based games, offering both hand-authored and procedurally generated games. Game-based environments also emphasize visual settings that may evolve independently. Crafter (Hafner, 2022) is a 2D open-world survival game that benchmarks deep exploration and long-horizon reasoning. Craftax (Matthews et al., 2024), built upon Crafter using JAX, introduces increased complexity and GPU-acceleration for open-ended RL. The modified Crafter variant by ELLM (Du et al., 2023) expands the action space and introduces distractor tasks. For multi-agent coordination, SMAC (Samvelyan et al., 2019) and SMAC-Hard (Deng et al., 2024b) provide StarCraft II-based benchmarks for cooperative decentralized control. SMAC-R1 (Deng et al., 2024b), Adaptive Command (Ma et al., 2025b) and TacticCraft (Ma et al., 2025a) further advance the performance of LLM agents in StarCraft II-style environments. Factorio (Hopkins et al., 2025) presents a dynamic, tick-based industrial simulation where agent inaction still alters the world state.
@@ 5.1.6 General-Purpose Environments
Some environments and benchmarks are designed for broad evaluation or to improve general agent capabilities. AgentGym (Xi et al., 2025) focuses on improving LLM agent generalization via instruction tuning and self-correction, operating on deterministic environments such as ALFWorld, BabyAI, and SciWorld. Agentbench (Liu et al., 2024b) serves as a broad evaluation framework, assessing LLMs as agents across a variety of distinct interactive environments, including SQL-based, game-based, and web-based scenarios. InternBootcamp (Li et al., 2025g) is a scalable framework integrating over 1000 verifiable reasoning tasks, spanning programming, logic puzzles, and games, with a standardized interface for RL training and automated task generation.